# Ambitions for theory in the physics of life

### v1

@@ 1.4 Agenda
This explosion of experimental developments obviously creates a need for new methods of data analysis. Indeed, as these approaches to high dimensional data collection have penetrated the mainstream of biology, even the biology community itself has emphasized the urgency of this need for mathematical analysis. But for physicists theory is more than data analysis. The most powerful analyses are grounded in theories, and one could make a strong case that all analysis methods embody some theoretical prejudice. [^15] In particle physics, for example, the signals from thousands of detector elements are reduced to a plot of the rate for some class of events vs some energy variable. An enormous amount of theoretical understanding is contained in the idea that this is the right thing to plot, even before we ask theory to predict what the plot shows. Something we do well in physics is to make these theoretical prejudices explicit. [^16] At the risk of being pedantic, consider the simple idea that high dimensional data—expression levels of hundreds of genes, electrical activity of thousands of neurons, and more—live in low dimensional spaces. This seems “theory free,” testable by standard methods for linear [117] or nonlinear [118] dimensionality reduction. But low dimensionality is a theoretical claim: what is the principle that limits the dimensionality of the dynamics? More subtly, to measure how well a low dimensional description works, we need a metric, and this metric is a theoretical claim about which variations are most relevant, or perhaps which variations are measured most reliably. Returning to the fly embryo, as we discussed at the outset any attempt to give a realistic quantitative description immediately leads us into a dense forest of parameters. Do we have clues about what principle(s) might cut through this complexity?
@@
The system is extraordinarily precise. It has been thought for decades that just $\sim 3\,{\rm hrs}$ after the egg is laid, every cell along the anterior–posterior axis knows its fate [119]. Apparently the genetic network schematized in Fig 2 carries enough information to do this. Since cell fates are tied to positions, and there are fewer than 100 rows of cells along the anterior–posterior axis, this means that the concentrations of just a handful of molecules specify position with $\sim 1\%$ accuracy. This can be made explicit, e.g. by measuring the reproducibility of the pair rule stripe positions as in Fig 8 [120, 121].
@@
Figure 8: Pair rule stripes have positional noise $\sigma_{x}/L\sim 0.01$. (A) Positions of the seven pair rule stripes for the gene shown at bottom in Fig 2 drift with time during the fourteenth cycle of nuclear divisions, illustrated here for 100+ embryos. (B) Once corrected for drift, the variance of stripe positions is very small. Results for seven stripes from three different pair rule genes (Fig 1, bottom); error bars are standard errors of the mean. Dashed line is a positional noise equal to one percent of the embryo length. Redrawn from Refs [120, 121].
@@
The concentrations of relevant molecules are low. Almost all the molecules in the network of Fig 2 are transcription factors—proteins that bind to specific sites along DNA and regulate the expression of other genes, in this case other genes in the network. Again it has been known for decades that transcription factors function at concentrations measured in nanoMolar [122, 123, 124], and there is no reason to doubt that this is true of the relevant molecules in the embryo [125]. Cell nuclei have dimensions measured in microns, and $1\,{\rm nM}=0.6\,{\rm molecules}/\mu{\rm m}^{3}$, so the absolute numbers of molecules can’t be very large.
@@
These two facts (here somewhat stylized) might be in conflict—at low concentrations things are noisy (because of physics not biology), and it is hard for molecules to convey much information. On the other hand there might be a principle that reconciles the conflict: parameters of the relevant networks have been selected to convey as much information as possible from these physically limited signals. This principle that will be the focus of what follows.
@@
Optimization comes along with an aesthetic that you might or might not find appealing. But what matters is that optimization principles make quantitative predictions that can be tested in modern experiments. [^17] See also Chapter 3 in Ref [57]. In many cases, as we will see, these predictions are essentially parameter free and accomplish the goal of circumventing highly parameterized models. Importantly, the claim that information flow is being optimized can be tested by measuring the information flow itself, in bits or as an effective noise level against which independently measured signals must be compared in order for the system to function. If we go back to the picture of information flow from maternal inputs through the gap genes to the pair rule stripes, we can identify three distinct opportunities for optimization.
Optimal decoding. Information about position is encoded in the concentrations of the four gap gene products. Given the (measured) noise levels in these concentrations, there is an optimal strategy for decoding this information, extracting an estimate of position that is as accurate as possible. Does this optimal precision match the observed $\sim 1\%$ precision of the pair rule stripes? Can we test whether the embryo really implements something functionally equivalent to the optimal decoding algorithm?
@@
Network architecture. Finally, we can go back to the 50+ parameters of gap gene network and ask if these parameters have been set to optimize information transmission. This is a hard problem, and we have tried to make progress by breaking off small pieces. Recently there has been a major step forward, optimizing the whole gap gene network over a class of models that is almost (rich) enough to include the real network. We are coming close to deriving the properties of this network from a general physical principle, with no free parameters.
@@ 2 Optimal decoding
If we think that the genetic regulatory networks in the embryo have been selected to transmit as much positional information as possible while using a limited number of molecules, it would be very odd if the cells then used this information inefficiently. So it is reasonable to ask how the embryo could use the concentrations of the gap proteins to draw the most reliable inferences about position, and whether we can find signature of this optimization. We’ll need to build up some tools to answer these questions.
@@ 2.1 A warmup exercise
In the case where both the signal $x$ and the noise $\eta$ are Gaussian, as above, then these two different definitions of the optimal estimate agree. More generally if different but plausible definitions of “best” lead to significantly different estimator, it probably is a sign that $P(x|y)$ has a complicated structure, such as multiple peaks, so that inference is not just noisy but genuinely ambiguous.
@@ 2.2 Counting photons and estimating motion
There is much to discuss here, much of it now classical but still some open questions. I want to focus on one part of the retinal circuitry problem.
@@
Although $y_{\rm i}$ is on average equal to $n_{\rm i}$ the best estimate of photon count is not the receptor output itself. Following argument above, we have
@@
Equation (24) says that, in the presence of motion, the spatial and temporal derivatives should be proportional to one another. A gentler statement is that these derivative are correlated, and the strength of this correlation should be related to the movement velocity. This suggests that we might be able to estimate
@@
Figure 10: Estimation of visual motion from local derivatives [141]. The optimal estimator in Eq (26) is computed from data collected by a special purpose camera during a walk in the woods. Black line indicates the predictions of Eq (24) for $v_{\rm deriv}=70\,^{\circ}/{\rm s}$. White square encloses $90\%$ of the data.
@@ 2.3 Concentration measurements, revisited
Figure 11: Expression levels of the four gap genes as a function of position along the long (anterior–posterior) axis of the embryo. Bottom panel shows the means $\langle g_{\rm i}(x)\rangle$ (blue lines) and standard deviations $\sqrt{\langle[\delta g_{\rm i}(x)]^{2}\rangle}$ (cyan shading). Top panels shows examples of the covariance matrix $C_{\rm ij}(x)=\langle\delta g_{\rm i}(x)\delta g_{\rm j}(x)\rangle$ at three positions, $x/L=0.3,0.5,$ and $0.7$. Data from Ref [10].
Figure 11 shows the spatial profiles of the gap gene expression levels along the long (anterior–posterior) axis of the embryo. [^21] I tried to give this course without mentioning the names of these (and other) molecules, because I don’t think it matters. But, to connect with the literature, the names are (1) Hunchback, (2) Krüppel, (3) Giant, and (4) Knirps. These data are extracted from the fluorescent antibody staining experiments discussed in §1.3. We will refer to the concentration of each molecule at position $x$ as $g_{\rm i}(x)$, with the index i=1, 2, 3, ${\rm i}=1,\,2,\,3,\,4$. Solid lines show the mean concentrations $\langle g_{\rm i}(x)\rangle$, with cyan shading to indicate the standard deviation of fluctuations around this mean. [^22] Recall from Fig 4 that we choose units such that $\langle g_{\rm i}(x/L)\rangle$ runs between 0 and 1 for each gene. An important qualitative observation is that these fluctuations in fact are quite small. These data sets are large enough that can estimate, reliably, the $4\times 4$ covariance matrix of fluctuations at each point,
@@
Armed with this distribution we can compute many things.
In particular we could think about making a single estimator $x_{\rm est}\left(\{g_{\rm i}\}\right)$. As above, this could be the optimal estimator in the least–square sense, which is the conditional mean
@@
If the distribution $P\left(x|\{g_{\rm i}\}\right)$ has a single sharp peak in the neighborhood of $x_{\rm est}^{(2)}\left(\{g_{\rm i}\}\right)$, then these two estimators will be very close to another, and to any other reasonable estimator, e.g. the one that minimize the mean absolute error (the $L_{1}$ estimator). On the other hand, if there is genuine ambiguity, so that $P\left(x|\{g_{\rm i}\}\right)$ has more than one peak, or if the estimation is very uncertain, so that $P\left(x|\{g_{\rm i}\}\right)$ is extremely broad, then no single estimator really captures what a cell “knows” based on the expression levels $\{g_{\rm i}\}$. At the start, it is not obvious that cells won’t be in one of these ambiguous or uncertain situations, so we would like to keep all the available information. This requires us to visualize $P\left(x|\{g_{\rm i}\}\right)$ more directly.
@@
Figure 12: Decoding position from a single gene. (A) Top panel is an optical section through the midsagittal plane of a Drosophila embryo with immunofluorescence labeling for the protein in encoded by a single gap gene, corresponding to $g_{2}$ in Fig 11 (scale bar $100\,\mu{\rm m}$). Bottom panel shows the normalized $g_{2}(x)$ for this particular embryo (blue) and a gray–level visualization of $P(g_{2}|x)$; $P(g_{2})$ is at left. (B) The distribution $P(x|g_{2})$, shown in gray levels (bottom) and slices at fixed $g_{2}$ (top).(C) The average decoding map $P(x^{*}|x)$ from Eq (37). Top panel schematizes the combination of ambiguity and uncertainty. Redrawn from Ref [10].
@@
Figure 13: Decoding position from an increasing number of gap genes reduces uncertainty and eliminates ambiguity. Redrawn from Ref [10].
@@
We recall that the pair rule stripes are positioned with $\sim 1\%$ accuracy, and similarly the location of the cephalic furrow is reproducible from embryo to embryo with $\sim 1\%$ precision. A possible conclusion is that embryos “read out” the information carried by local gap gene concentrations and use this to guide subsequent events. This read out is optimal, and sets the precision of the body plan. Certainly what we see is consistent with this conclusion, but the evidence is not unambiguous.
@@
Figure 14: Gap gene concentration profiles change in mutant flies missing one of the maternal inputs. At left are the profiles in wild type flies, the same information as in Fig 11 but with the four genes represented by different colors. At right are the results of the same measurements done in flies that are missing the terminal inputs T. Redrawn from Ref [10]. It is essential that although concentrations are normalized (maximum mean concentration is one in the wild type), the normalization is the same in both panels so that concentrations can be compared meaningfully.
Figure 15 shows what happens when we try to decode the patterns of gap gene expression found in the mutants of Fig 14. The idea is that we take the measured $\{g_{\rm i}\}$ in the mutant and pass it through the function $P(x^{*}|\{g_{\rm i}\})$, which was constructed from data the normal (wild—type) embryos, and then we average as before to get a new map $P_{\rm mut}(x^{*}|x)$. By analogy with Eq (37) we can write
@@
Figure 15: Optimal decoding predicts pair rule stipe shifts in mutant embryos. We use the distribution $P(x^{*}|\{g_{\rm i}\})$ constructed from data on wild type flies to interpret the gap gene expression signals in flies missing terminal inputs T (left panel in Fig 14). Results are shown as the average decoding map $P_{\rm mut}(x^{*}|x)$ from Eq (38); density in grey scale as shown bottom left. At left is the expression of Eve, one of the pair rule genes; individual embryos in grey, mean in black, illustrating the seven stripes at positions $x_{\rm i}$. When $P_{\rm mut}(x^{*}=x_{\rm i}|x)$ is large we expect a stripe at (real) position $x$ in the mutant, as confirmed at the bottom. Markers in the main figure summarize these results, and those for two other pair rule genes. Redrawn from Ref [10].
We can test the predicted $P_{\rm mut}(x^{*}|x)$ with a simple idea. If the only information that cells have about position is their estimate $x^{*}$, then pair rule stripes should be at locations $x^{*}=x_{\rm s}$ not $x=x_{\rm s}$ as in the wild–type embryo, for each stripe $\rm s$; this is the construction shown by dashed lines in Fig 15 for the gene eve. [^23] This is one place where I’ll use names. Eve is short for “even–skipped,” which gives a sense for what goes wrong in the structure of the mutant embryo. The convention is that genes are named in lower case italics, while the protein products are described with capitalized Roman text. Also, beware that biochemists name molecules after what they do, while geneticists name molecules after what goes wrong when they are mutated. We can test all of the stripes, for each of several pair rule genes, and we see that the results track along the ridge of $P_{\rm mut}(x^{*}|x)$, with high precision.
@@
  • We can analyze maps $P^{\alpha}_{\rm mut}(x^{*}|x)$ constructed from data on individual mutant embryos (before averaging), and in some cases the density is sufficiently variable at $x^{*}=x_{s}$ that we predict stripes to be present in some but not all of the embryos. This variability never happens in wild–type embryos but it happens in the mutants, where we predict it. [^24] A limitation of the experiments is that they measures all the gap genes or the pair rule genes, but not in the same embryo. Thus we know that pair rule stripes are variable where we predict them to be variable based on measurements of the gap genes in many mutant embryos, but we don’t know if the presence or absence of stripe is connected deterministically to the gap gene profiles in single embryos, as we predict.
  • In mutants where the anterior input is deleted, the most likely $x^{*}$ is a non–monotonic function of $x$. Most of the eve stripes are predicted to be missing, but the seventh stripe is predicted to be duplicated, and this happens the predicted location. Details of the underlying molecular biology show that this really is a duplicate of stripe seven and not a shifted version of a more anterior stripe.
@@
We have constructed the optimal decoder $P(x^{*}|\{g_{\rm i}\})$ from a measurements in a small window of time during the fourteenth nuclear cycle. This window is chosen to surround the point at which positional information is maximized, and is as narrow as possible while still leaving a reasonable number of samples. But gap gene expression profiles vary slowly throughout cycle fourteen. If the embryo implements the optimal decoding, tuned to the time of maximal positional information, then pair rule stripes are predicted to evolve with time as well. This effect has been known for a long time, and subject to multiple interpretations. We were surprised to find that these details—shifts of $\Delta x(t)/L$ corresponding to just a few percent over half an hour—are predicted correctly as well.

### v2

@@ 1.4 Agenda
This explosion of experimental developments obviously creates a need for new methods of data analysis. Indeed, as these approaches to high dimensional data collection have penetrated the mainstream of biology, the biology community itself has emphasized the urgency of this need for mathematical analysis. But for physicists theory is more than data analysis. The most powerful analyses are grounded in theories, and there is a strong case that all analysis methods embody some theoretical prejudice. [^15] In particle physics, for example, the signals from thousands of detector elements are reduced to a plot of the rate for some class of events vs some energy variable. An enormous amount of theoretical understanding is contained in the idea that this is the right thing to plot, even before we ask theory to predict what the plot shows. Something we do well in physics is to make these theoretical prejudices explicit. [^16] At the risk of being pedantic, consider the simple idea that high dimensional data—expression levels of hundreds of genes, electrical activity of thousands of neurons, and more—live in low dimensional spaces. This seems “theory free,” testable by standard methods for linear [117] or nonlinear [118] dimensionality reduction. But low dimensionality is a theoretical claim: what is the principle that limits the dimensionality of the dynamics? More subtly, to measure how well a low dimensional description works, we need a metric, and this metric is a theoretical claim about which variations are most relevant, or perhaps which variations are measured most reliably. Returning to the fly embryo, as we discussed at the outset any attempt to give a realistic quantitative description immediately leads into a dense forest of parameters. Do we have clues about what principle(s) might cut through this complexity?
@@
The system is extraordinarily precise. It has been thought for decades that just $\sim 3\,{\rm hrs}$ after the egg is laid, every cell along the anterior–posterior axis knows its fate [119]. Apparently the genetic network schematized in Fig 2 carries enough information to do this. Since cell fates are tied to positions, and there are fewer than 100 rows of cells along the anterior–posterior axis, this means that the concentrations of just a handful of molecules specify position with $\sim 1\%$ accuracy. This can be made explicit, e.g. by measuring the reproducibility of the pair–rule stripe positions as in Fig 8 [120, 121].
@@
Figure 8: Pair–rule stripes have positional noise $\sigma_{x}/L\sim 0.01$. (A) Positions of the seven pair–rule stripes for the gene shown at bottom in Fig 2 drift with time during the fourteenth cycle of nuclear divisions, illustrated here for 100+ embryos. (B) Once corrected for drift, the variance of stripe positions is very small. Results for seven stripes from three different pair–rule genes (Fig 1, bottom); error bars are standard errors of the mean. Dashed line is a positional noise equal to one percent of the embryo length. Redrawn from Refs [120, 121], with thanks to H Casademunt, JO Dubuis, T Gregor, L McGough, M Nikolić, MD Petkova, G Tkačik, and EF Wieschaus.
@@
The concentrations of relevant molecules are low. Almost all the molecules in the network of Fig 2 are transcription factors—proteins that bind to specific sites along DNA and regulate the expression of other genes, in this case other genes in the network. Again it has been known for decades that transcription factors function at concentrations measured in nanoMolar [122, 123, 124], and there is no reason to doubt that this is true of the relevant molecules in the embryo [125]. Cell nuclei have dimensions measured in microns, and $1\,{\rm nM}=0.6\,{\rm molecules}/\mu{\rm m}^{3}$, so even the absolute numbers of molecules can’t be very large.
@@
These two facts (here somewhat stylized) might be in conflict—at low concentrations things are noisy (because of physics not biology), and it is hard for molecules to convey much information. On the other hand there might be a principle that reconciles the conflict: parameters of the relevant networks have been selected to convey as much information as possible from these physically limited signals. This principle will be the focus of what follows.
@@
Optimization comes along with an aesthetic that you might or might not find appealing. But what matters is that optimization principles make quantitative predictions that can be tested in modern experiments. [^17] See also Chapter 3 in Ref [57]. In many cases, as we will see, these predictions are essentially parameter free and accomplish the goal of circumventing highly parameterized models. Importantly, the claim that information flow is being optimized can be tested by measuring the information flow itself, in bits or as an effective noise level against which independently measured signals must be compared in order for the system to function. If we go back to the picture of information flow from maternal inputs through the gap genes to the pair–rule stripes, we can identify three distinct opportunities for optimization.
Optimal decoding. Information about position is encoded in the concentrations of the four gap gene products. Given the (measured) noise levels in these concentrations, there is an optimal strategy for decoding this information, extracting an estimate of position that is as accurate as possible. Does this optimal precision match the observed $\sim 1\%$ precision of the pair–rule stripes? Can we test whether the embryo really implements something functionally equivalent to the optimal decoding algorithm?
@@
Network architecture. Finally, we can go back to the 50+ parameters of the gap gene network and ask if these parameters have been set to optimize information transmission. This is a hard problem, and we have tried to make progress by breaking off small pieces. Recently there has been a major step forward, optimizing the whole gap gene network over a class of models that is almost (rich) enough to include the real network. We are coming close to deriving the properties of this network from a general physical principle, with no free parameters.
@@ 2 Optimal decoding
If we think that the genetic regulatory networks in the embryo have been selected to transmit as much positional information as possible while using a limited number of molecules, it would be very odd if the cells then used this information inefficiently. So it is reasonable to ask how the embryo could use the concentrations of the gap proteins to draw the most reliable inferences about position, and whether we can find signatures of this optimization. We’ll need to build up some tools to answer these questions.
@@ 2.1 A warmup exercise
In the case where both the signal $x$ and the noise $\eta$ are Gaussian, as above, then these two different definitions of the optimal estimate agree. More generally if different but plausible definitions of “best” lead to significantly different estimators, it probably is a sign that $P(x|y)$ has a complicated structure, such as multiple peaks, so that inference is not just noisy but genuinely ambiguous.
@@ 2.2 Counting photons and estimating motion
There is much to discuss here, much of it now classical but still some questions remain open. I want to focus on one part of the retinal circuitry problem.
@@
Although $y_{\rm i}$ is on average equal to $n_{\rm i}$ the best estimate of photon count is not the receptor output itself. Following the argument above, we have
@@
Equation (24) says that, in the presence of motion, the spatial and temporal derivatives should be proportional to one another. A gentler statement is that these derivatives are correlated, and the strength of this correlation should be related to the movement velocity. This suggests that we might be able to estimate
@@
Figure 10: Estimation of visual motion from local derivatives. The optimal estimator in Eq (26) is computed from data collected by a special purpose camera during a walk in the woods. Black line indicates the predictions of Eq (24) for $v_{\rm deriv}=70\,^{\circ}/{\rm s}$. White square encloses $90\%$ of the data. Redrawn from Ref [141], with thanks to S Sinha and RR de Ruyter van Steveninck.
@@ 2.3 Concentration measurements, revisited
Figure 11: Expression levels of the four gap genes as a function of position along the long (anterior–posterior) axis of the embryo. Bottom panels show the means $\langle g_{\rm i}(x)\rangle$ (blue lines) and standard deviations $\sqrt{\langle[\delta g_{\rm i}(x)]^{2}\rangle}$ (cyan shading). Top panels show examples of the covariance matrix $C_{\rm ij}(x)=\langle\delta g_{\rm i}(x)\delta g_{\rm j}(x)\rangle$ at three positions, $x/L=0.3,0.5,$ and $0.7$. Data from Ref [10], with thanks to T Gregor, MD Petkova, G Tkačik, and EF Wieschaus.
Figure 11 shows the spatial profiles of the gap gene expression levels along the long (anterior–posterior) axis of the embryo. [^21] I tried to give this course without mentioning the names of these (and other) molecules, because I don’t think it matters. But, to connect with the literature, the names are (1) Hunchback, (2) Krüppel, (3) Giant, and (4) Knirps. These data are extracted from the fluorescent antibody staining experiments discussed in §1.3. We will refer to the concentration of each molecule at position $x$ as $g_{\rm i}(x)$, with the index i=1, 2, 3, ${\rm i}=1,\,2,\,3,\,4$. Solid lines show the mean concentrations $\langle g_{\rm i}(x)\rangle$, with cyan shading to indicate the standard deviation of fluctuations around this mean. [^22] Recall from Fig 4 that we choose units such that $\langle g_{\rm i}(x/L)\rangle$ runs between 0 and 1 for each gene. An important qualitative observation is that these fluctuations in fact are quite small. These data sets are large enough that we can estimate, reliably, the $4\times 4$ covariance matrix of fluctuations at each point,
@@
From the distribution $P\left(x|\{g_{\rm i}\}\right)$ we can compute many things.
In particular it is tempting to think about constructing a single estimator $x_{\rm est}\left(\{g_{\rm i}\}\right)$. As above, this could be the optimal estimator in the least–square sense, the conditional mean
@@
If the distribution $P\left(x|\{g_{\rm i}\}\right)$ has a single sharp peak in the neighborhood of $x_{\rm est}^{(2)}\left(\{g_{\rm i}\}\right)$, then these two estimators will be very close to one another, and to any other reasonable estimator, e.g. the one that minimizes the mean absolute error (the $L_{1}$ estimator). On the other hand, if there is genuine ambiguity, so that $P\left(x|\{g_{\rm i}\}\right)$ has more than one peak, or if the estimation is very uncertain, so that $P\left(x|\{g_{\rm i}\}\right)$ is extremely broad, then no single estimator really captures what a cell “knows” based on the expression levels $\{g_{\rm i}\}$. At the start, it is not obvious that cells won’t be in one of these ambiguous or uncertain situations, so we would like to keep all the available information. This requires us to visualize $P\left(x|\{g_{\rm i}\}\right)$ more directly.
@@
Figure 12: Decoding position from a single gene. (A) Top panel is an optical section through the midsagittal plane of a Drosophila embryo with immunofluorescence labeling for the protein encoded by a single gap gene, corresponding to $g_{2}$ in Fig 11 (scale bar $100\,\mu{\rm m}$). Bottom panel shows the normalized $g_{2}(x)$ for this particular embryo (blue) and a gray–level visualization of $P(g_{2}|x)$; $P(g_{2})$ is at left. (B) The distribution $P(x|g_{2})$, shown in gray levels (bottom) and slices at fixed $g_{2}$ (top). (C) The average decoding map $P(x^{*}|x)$ from Eq (37). Top panel schematizes the combination of ambiguity and uncertainty. Reproduced from Ref [10], with permission.
@@
Figure 13: Decoding position from an increasing number of gap genes reduces uncertainty and eliminates ambiguity. Redrawn from Ref [10], with permission.
@@
We recall that the pair–rule stripes are positioned with $\sim 1\%$ accuracy, and similarly the location of the cephalic furrow is reproducible from embryo to embryo with $\sim 1\%$ precision. A possible conclusion is that embryos “read out” the information carried by local gap gene concentrations and use this to guide subsequent events. This read out is optimal, and sets the precision of the body plan. Certainly what we see is consistent with this conclusion, but the evidence is not unambiguous.
@@
Figure 14: Gap gene concentration profiles change in mutant flies missing one of the maternal inputs. At left are the profiles in wild type flies, as in Fig 11 but with the four genes represented by different colors. At right are the results of the same measurements done in flies that are missing the terminal inputs T. Reproduced from Ref [10], with permission. It is essential that although concentrations are normalized (maximum mean concentration is one in the wild type), the normalization is the same in both panels so that concentrations can be compared meaningfully.
Figure 15 shows what happens when we try to decode the patterns of gap gene expression found in the mutants of Fig 14. The idea is that we take the measured $\{g_{\rm i}\}$ in the mutant and pass it through the function $P(x^{*}|\{g_{\rm i}\})$, which was constructed from data taken in the normal (wild–type) embryos, and then we average as before to get a new map $P_{\rm mut}(x^{*}|x)$. By analogy with Eq (37) we can write
@@
Figure 15: Optimal decoding predicts pair–rule stipe shifts in mutant embryos. We use the distribution $P(x^{*}|\{g_{\rm i}\})$ constructed from data on wild type flies to interpret the gap gene expression signals in flies missing terminal inputs T (Fig 14). Results are shown as the average decoding map $P_{\rm mut}(x^{*}|x)$ from Eq (38); density in grey scale (bottom left). At left is the expression of Eve, one of the pair–rule genes; individual embryos in grey, mean in black, illustrating the seven stripes at positions $x_{\rm i}$. When $P_{\rm mut}(x^{*}=x_{\rm i}|x)$ is large we expect a stripe at (real) position $x$ in the mutant, as confirmed at the bottom. Markers summarize these results, and those for two other pair–rule genes. Reproduced from Ref [10], with permission.
We can test the predicted $P_{\rm mut}(x^{*}|x)$ with a simple idea. If the only information that cells have about position is their estimate $x^{*}$, then pair–rule stripes should be at locations $x^{*}=x_{\rm s}$ not $x=x_{\rm s}$ as in the wild–type embryo, for each stripe $\rm s$; this is the construction shown by dashed lines in Fig 15 for the gene eve. [^23] This is one place where I’ll use names. Eve is short for “even–skipped,” which gives a sense for what goes wrong in the structure of the mutant embryo. The convention is that genes are named in lower case italics, while the protein products are described with capitalized Roman text. Also, beware that biochemists name molecules after what they do, while geneticists name molecules after what goes wrong when they are mutated. We can test all of the stripes, for each of several pair–rule genes, and we see that the results track along the ridge of $P_{\rm mut}(x^{*}|x)$, with high precision.
@@
  • We can analyze maps $P^{\alpha}_{\rm mut}(x^{*}|x)$ constructed from data on individual mutant embryos (before averaging), and in some cases the density is sufficiently variable at $x^{*}=x_{s}$ that we predict stripes to be present in some but not all of the embryos. This variability never happens in wild–type embryos but it happens in the mutants, where we predict it. [^24] A limitation of the experiments is that they measure all the gap genes or the pair–rule genes, but not in the same embryo. Thus we know that pair–rule stripes are variable where we predict them to be variable based on measurements of the gap genes in many mutant embryos, but we don’t know if the presence or absence of stripes is connected deterministically to the gap gene profiles in single embryos, as we predict.
  • In mutants where the anterior input is deleted, the most likely $x^{*}$ is a non–monotonic function of $x$. Most of the eve stripes are predicted to be missing, but the seventh stripe is predicted to be duplicated, and this happens at the predicted location. Details of the underlying molecular biology show that this really is a duplicate of stripe seven and not a shifted version of a more anterior stripe.
@@
We have constructed the optimal decoder $P(x^{*}|\{g_{\rm i}\})$ from measurements in a small window of time during the fourteenth nuclear cycle. This window is chosen to surround the point at which positional information is maximized, and is as narrow as possible while still leaving a reasonable number of samples. But gap gene expression profiles vary slowly throughout cycle fourteen. If the embryo implements the optimal decoding, tuned to the time of maximal positional information, then pair–rule stripes are predicted to evolve with time as well. This effect has been known for a long time, and subject to multiple interpretations. We were surprised to find that these details—shifts of $\Delta x(t)/L$ corresponding to just a few percent over half an hour—are predicted correctly as well.